Find the change in kinetic energy by an elastic spring with spring constant of 100 N/cm, as it goes back to its equilibrium position, from a length of 10 cm to 15 cm.


a. -250 J


b. -1250 J


c. 0


d. -2500 J


1
Expert's answer
2022-03-13T18:48:42-0400

The potential energy stored in the compressed string is given as follows:


"E_p = \dfrac{kx^2}{2}"

where "x = 15cm-10cm = 5cm = 0.05m" is the contraction, "k = 100N\/cm = 10^4N\/m" is the spring constant.

As the spring returns in the equilibrium position, all this potential energy converts to the kinetic energy. Thus, the change in kinetic energy will be:


"\Delta K = \dfrac{kx^2}{2} = \dfrac{10^4N/m\cdot (0.05m)^2}{2} = 12.5J"

Answer. 12.5J.